OpenStudy (anonymous):

Through(-5,1) slope (- 6 over 5)

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

y = mx + b, m = slope = -6/5, x = -5, y = 1 Insert all known values above into above equation: 1 = (-6/5)(-5) + b Solve for b 1= 6 + b b = -5 Now that you know m and b, insert them back into slope-intercept equation: y = (-6/5)x - 5

OpenStudy (anonymous):

thank you again i may need some other help in a bit

OpenStudy (anonymous):

find the slope of a line perpendicular to each given line ...... y = - 4 over 5 x -1

OpenStudy (mathstudent55):

I'll do this one, but then I need to go. The slope of perpendicular lines are negative reciprocals. Reciprocals are numbers that multiply to 1. Negative reciprocals are numbers that multiply to -1. This explanation may seem long, but the concept is simple to do. How do you find the reciprocal of a number? First, write the number as a fraction. Then, flip the fraction. That's the reciprocal. Since for the slope of a perpendicular line you need a negative reciprocal, then first write the slope of the given line as a fraction, then flip the fraction and change the sign. There's your negative reciprocal. For example, a line has slope 5/6. What isd the slope of a perpendicular? Flip 5/6 and get 6/5. Then change the sign. Since 6/5 is positive, make it negative. Answer: -6/5 Another example: A line has slope -2. What is the slope of a perpendicular to this line? First write -2 as a fraction: -2/1. Now flip it: -1/2. Now change the sign. Answer: 1/2

OpenStudy (mathstudent55):

Now let's look at your problem. Find the slope of a line perpendicular to each given line ...... y = - 4 over 5 x -1 y = (-4/5)x - 1 This line is already in slope-intercept form, y = mx + b. The slope is m, so in this case the slope is -(4/5) To get the slope of a perpendicular, you need the negative reciprocal of -(4/5). So flip it and change the sign: 5/4